# DPA: Dual Prototypes Alignment for Unsupervised Adaptation of Vision-Language Models

### v1

@@ Abstract
Vision-language models (VLMs), e.g., CLIP, have shown remarkable potential in zero-shot image classification. However, adapting these models to new domains remains challenging, especially in unsupervised settings where labelled data is unavailable. Recent research has proposed pseudo-labelling approaches to adapt CLIP in an unsupervised manner using unlabelled target data. Nonetheless, these methods struggle due to noisy pseudo-labels resulting from the misalignment between CLIP’s visual and textual representations. This study introduces DPA, an unsupervised domain adaptation method for VLMs. DPA introduces the concept of dual prototypes, acting as distinct classifiers, along with the convex combination of their outputs, thereby leading to accurate pseudo-label construction. Next, it ranks pseudo-labels to facilitate robust self-training, particularly during early training. Finally, it addresses visual-textual misalignment by aligning textual prototypes with image prototypes to further improve the adaptation performance. Experiments on 13 downstream vision tasks demonstrate that DPA significantly outperforms zero-shot CLIP and the state-of-the-art unsupervised adaptation baselines.
@@ 1 Introduction
Context and background: Vision-language models (VLMs) [1, 2, 3, 4] have shown promising potential in zero-shot image classification. One notable example is CLIP [1]. Despite the impressive zero-shot capabilities of CLIP, its performance can be impacted by the discrepancy between the pretraining image-text pairs and the downstream task images [1, 5]. To address this limitation, several studies have attempted to enhance CLIP transfer performance on downstream tasks by leveraging limited labelled samples from the target domains [6, 7, 8, 9, 10]. However, in many practical applications, such as security and medical diagnostics, collecting labelled samples can be particularly challenging due to high labelling costs and data privacy concerns. In such contexts, unsupervised learning presents a promising alternative. Several contributions for adapting CLIP to a target domain using an unlabelled dataset have been introduced recently [11, 12, 13, 14, 15, 16]. A common approach involves leveraging CLIP to generate pseudo-labels, which are then used to fine-tune CLIP in an unsupervised manner. However, this method encounters significant challenges due to the noisy pseudo-labels generated by CLIP. Despite efforts by recent methods for the unsupervised adaptation of CLIP, a significant modality gap persists between the text and vision representations [16, 17], leading to inaccurate pseudo-labels causing confirmation bias during adaptation.
Motivation: Recent studies highlight a significant factor behind the performance issues of CLIP in unsupervised adaptation scenarios: the visual domain gap between the source images used to train CLIP and target images typically occurs when the target samples originate from an uncommon domain [16]. As shown in Figure 1-(a), the t-SNE projection of visual and textual embeddings on the EuroSAT dataset [18] reveals a significant misalignment, resulting in misclassifications. Contemporary methods (e.g., [16]) attempt to address this misalignment by learning a projection space to mitigate the issue and using label propagation [19] to improve pseudo-labels. However, they could be limited as they are either computationally expensive and/or show suboptimal performance in an inductive setting. Further, on datasets with a large number of classes, they struggle to perform well, necessitating turning off label propagation and using model predictions. Effectively addressing the challenges of unsupervised adaptation of CLIP in an inductive setting is crucial for improving CLIP’s performance in unsupervised adaptation scenarios.
@@
Our proposal: To address these challenges, we propose DPA, a novel unsupervised adaptation method for VLMs (Figure 2). DPA aims at addressing the domain gap between visual and textual representations in downstream tasks. DPA introduces the idea of dual prototypes, namely image and textual prototypes, which act as distinct classifiers and their outputs are fused via convex combination towards generating accurate pseudo-labels. There are two main reasons for introducing dual prototypes. Firstly, the misalignment between the image representation and its textual representation in zero-shot CLIP, due to domain shift, often leads to inaccurate pseudo-labels (see Figure 1-(a)). Secondly, image prototypes, which tend to be less affected by noise, as illustrated in Figure 1, are generally closer to the true image representation than textual prototypes. Additionally, we tackle the challenge of misalignment between visual and textual embeddings by aligning textual prototypes with image prototypes. This alignment process further enhances the performance of unsupervised CLIP adaptation.
Contributions: 1) We propose DPA, a novel unsupervised domain adaptation framework for VLMs which allows adapting these powerful models to new domains without requiring labelled data from the target distribution. 2) We introduce a novel approach for generating accurate pseudo-labels by leveraging two distinct prototypes and fusing their outputs via a convex combination. We also propose to rank pseudo-labels in the classification loss to mitigate their noise, especially during early training. Moreover, we propose to tackle the visual-textual misalignment by aligning textual prototypes with image prototypes. 3) Experiments on 13 downstream vision tasks demonstrate consistent and significant performance enhancements over zero-shot CLIP and the state-of-the-art baselines.
@@ 2 Related Work
Large-scale Vision-Language (VL) Models: Among other vision-language models [20, 21, 2, 1, 22, 3, 23], CLIP [1] stands out as a pioneering example, aligning visual and textual features via a contrastive objective on a vast web-crawled collection of image-text pairs. This alignment empowers CLIP to generalize to diverse downstream classification tasks effectively. Building on CLIP’s success, subsequent research has explored methods for efficiently transferring the pre-trained model to handle diverse downstream tasks with limited labelled target data. These approaches often leverage vision-specific adapters [24, 6, 25, 26], language-specific adapters [7, 8, 10], or both [9]. However, these techniques require a minimum number of labelled samples, posing challenges due to data scarcity, annotation expenses, privacy issues, and practical limitations. Recent advancements have seen the emergence of unsupervised adaptation techniques, focusing on tailoring CLIP to target tasks using unlabelled datasets [14, 13, 16, 12, 11, 15]. However, these methods still depend on additional supervision signals, such as fine-tuning classifiers with a large language model like GPT-3, as seen in LaFTer [12]. They also often require significant computational resources, exemplified by MUST [11]. ReCLIP [16] tackle misaligned embeddings through source-free domain learning, utilizing a projection space to align these embeddings and employing pseudo-labels for self-training. Despite progress, there remains significant potential for improving the performance of unsupervised adaptation techniques across various image classification benchmarks. In contrast, our work introduces a fully label-free approach for adapting VLMs to a target task by generating more accurate pseudo-labels grounded on image and text prototypes.
Prototype-based Learning: Prototype-based learning uses a fixed set of distinctive prototypes representing the data, and then the learning is achieved by comparing the test samples directly with these prototypes. Prototype-based learning has been extensively studied in various contexts, including few-shot classification [27, 28], unsupervised learning [29, 30] and supervised classification [31, 32]. Our method innovatively incorporates two types of prototypes, image prototypes and textual prototypes, to harness both prototypes for ultimately generating accurate pseudo-labels and effectively adapting CLIP to the target dataset.
Pseudo-labelling: The success of pseudo-labelling has been demonstrated in various domains, including vision [33, 34] and vision-language tasks [11, 12, 14, 16]. The pseudo-labelling pipeline of DPA leverages consistency regularization [35] to produce consistent predictions across different data augmentations, thereby enhancing the accuracy of pseudo-labels.
@@ 3 Methodology
This study addresses the task of unsupervised adaptation of vision-language models (i.e., CLIP) for image classification. Following are the settings for this task:
  • Pre-trained CLIP consists of visual encoder $E_{v}$ and textual encoder $E_{t}$.
  • Target dataset $\mathcal{D}_{t}=\left\{\mathcal{X}_{t}\right\}$ consists of unlabelled images $\left\{x_{i}\right\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}_{t}$.
  • Unique class names $\mathcal{C}$ =$=\{c_{j}\}_{j=1}^{C}$for the unlabelled target dataset.
We use a pre-trained CLIP to process the unlabelled target data $\mathcal{D}_{t}$. Initially, the source model assigns a pseudo-label to each unlabelled target image $x_{i}$. During the adaptation phase, we employ two prototypes to generate pseudo-labels for the adaptation of CLIP to the target domain.
@@ 3.1 Zero-shot Visual Classification in CLIP
where $f_{i}^{s}=E_{v}(x_{i}^{s})$ and $z_{i}=E_{t}(\mathbf{t}^{s}_{i})$, with both $f_{i}^{s}\in\mathbb{R}^{1\times d}$ and $z_{i}\in\mathbb{R}^{1\times d}$. The parameter $\tau$ represents a learned temperature, while $m$ indicates the mini-batch size. For inference, given an unlabelled target image $x$ and a list of class names $\mathcal{C}$, we leverage the pre-trained textual encoder $E_{t}$ to generate text embeddings for each class. Specifically, a series of well-crafted prompts are fed into $E_{t}$, producing a set of text embeddings for each class, denoted as $\mathbf{Z}$ where $\mathbf{Z}\in\mathbb{R}^{C\times d}$. The zero-shot prediction $\hat{y}$ is then obtained as follows:
$\hat{y}=\operatornamewithlimits{argmax}_{c}(f\cdot\mathbf{Z}^{T})$ (2)
@@ 3.2 Proposed Framework (DPA)
Motivation: While zero-shot CLIP shows impressive performance across diverse domains, it still exhibits notable limitations compared to models adapted under supervised conditions due to the misalignment between image and corresponding textual representations [1]. As depicted in Figure 1-(a), this discrepancy is evident in CLIP when applied to uncommon domains. Such disparities can lead to inaccurate pseudo-labels, especially when utilizing zero-shot CLIP to generate pseudo-labels for the unsupervised adaptation [16]. Our objective is to adapt CLIP to the target domain using its unlabelled samples by bridging the gap between image representations and their textual representations, all without dependence on ground-truth labels. Simultaneously, we seek to achieve efficient parameter-based unsupervised adaptation for CLIP. Our objective relies on two key components: (1) ensuring the accuracy of the pseudo-labels generated for the unlabelled samples, and (2) aligning textual representations with their corresponding visual representations. To generate the pseudo-labels for adapting CLIP to the target domain, we introduce dual prototypes which act as two different classifiers and fuse their complementary outputs via convex combination. Furthermore, we propose visual-textual prototype alignment towards further improving the adaptation performance.
@@
Image Prototypes Construction: To construct the image prototypes, CLIP’s image encoder processes a weakly-augmented unlabelled image $\alpha(x)$ and generates a feature representation $f=E_{v}(\alpha(x))$, where $f\in\mathbb{R}^{1\times d}$. Without class labels, we utilize Equation 2 to generate pseudo-labels $\hat{y}$ from zero-shot CLIP. To obtain the image prototype $\mathbf{P}_{j}$ for a class $c_{j}$, we average the feature representation $f$ of the unlabelled images assigned to class $c_{j}$ as follows:
$\mathbf{P}_{j}=\frac{1}{N_{j}}\sum_{i=1}^{N}\mathbb{I}_{\hat{y_{i}}=j}f_{i}$ (4)
where $N_{j}=\sum_{i=1}^{N}\mathbb{I}_{\hat{y}_{i}=j}$. For the set of $C$ classes, we define the image prototypes as $\mathbf{P}\in\mathbb{R}^{C\times d}$. The image prototypes function as a non-parametric classifier. Given that the pseudo-labels generated by zero-shot CLIP are initially noisy, our approach, inspired by previous works [36, 37, 38], employs a memory bank $\mathbf{MB}$. This memory bank facilitates the incremental update of image prototypes in a non-parametric way throughout the training process. Specifically, $\mathbf{MB}$ stores the feature representations of all unlabelled target samples along with their corresponding pseudo-labels: $\mathbf{MB}=\{(f_{i},\hat{y}_{i})\}_{i=1}^{N}$. As the image prototypes are non-parametric, we iteratively refine them using the memory bank as follows: (1) We initially calculate the image prototypes using Equation 4, utilizing the image features and the pseudo-labels generated from zero-shot CLIP across the entire target dataset. (2) During training, we continuously update the memory bank by storing the image representations alongside their corresponding pseudo-labels. (3) At the end of each epoch, we update the prototypes $\mathbf{P}_{j}$ using Equation 4, incorporating the image representations and their corresponding pseudo-labels from $\mathbf{MB}$. (4) We repeat steps (2) and (3) until the completion of the training process. DPA maximizes the utilization of all available training samples to enhance the image prototypes. This iterative process ensures that each prototype continuously evolves, progressively assimilating relevant knowledge from individual training samples throughout the model training phase, while mitigating the error propagation commonly associated with pseudo-labels.
Pseudo-labels Generation: Both image and textual prototypes serve as distinct classifiers for the unlabelled target dataset. We generate a pseudo-label (PL) for the weakly-augmented unlabelled target image $\alpha(x)$ from these two prototypes. Initially, we determine the similarity between the image features and the textual prototypes as follows:
@@
Following [39], we perform distribution alignment (DA) to prevent the model’s prediction from collapsing onto specific classes. Here, $DA(p_{t})=p_{t}/\bar{p}_{t}$, where $\bar{p}_{t}$ represents a running average of $p_{t}$ during training, and $\beta$ is a hyperparameter. Finally, the PL is obtained as $\hat{y}=\operatornamewithlimits{argmax}_{c}\hat{p}$. This process of generating pseudo-labels improves the accuracy of the PL during training, indicating that these PLs may initially be noisy (see Figure 3). Therefore, assigning uniform weights to all pseudo-labels could hinder the adaptation process. To address this challenge, we propose adjusting the classification loss weight for the pseudo-label $\hat{y}$ based on its similarity to the textual and image prototypes, as follows:
@@
$\mathcal{L}_{st}=-\;\mathbb{E}_{x\in\mathcal{X}_{t}}\;\left[w_{x}\cdot\sum_{j=1}^{C}\mathbb{I}_{\hat{y}=j}log\,\left(p_{\mathcal{A}(x)}\right)\right],$ (9)
where $w_{x}$ represents the weight calculated from Equation 8, and $p_{\mathcal{A}(x)}=E_{v}(\mathcal{A}(x))\cdot\mathbf{Z}^{T}$ represents the probabilistic output for the strongly-augmented image $\mathcal{A}(x)$. To alleviate the confirmation bias induced by CLIP [40], we adopt “fairness” regularization as suggested in [11] as follows:
$\mathcal{L}_{reg}=-\frac{1}{C}\sum_{j=1}^{C}\log\left(\bar{p}_{\mathcal{A}(x),j}\right),$ (10)
where $\bar{p}_{\mathcal{A}(x)}$ represents the model’s average prediction from the strongly augmented images across the batch. By incorporating this fairness regularization, we aim to encourage the model to make more uniform predictions across classes, reducing the tendency to overfit the PLs and promoting a more balanced adaptation to the target domain. Equations 9 and 10 are used to update both the image encoder and the textual prototypes.
Prototypes Alignment: During training, the textual prototypes undergo continual updates to align with the associated image representation, leveraging pseudo-labels using Equation 9 and 10. Conversely, the refinement of the image prototypes occurs through averaging image features, as detailed in Section 3.2. We exploit the image prototypes to provide additional supervision for updating textual prototypes beyond pseudo-label alone. These image prototypes, less influenced by erroneous pseudo-labels and closer to image features (as illustrated in Figure 1), offer extra guidance for refining textual prototypes. Our primary objective is to optimize the relationship between image prototypes $\mathbf{P}$ and textual prototypes $\mathbf{Z}$ by maximizing the (cosine) similarity between these feature sets. To accomplish this optimization, we employ the InfoNCE loss [41] to directly assess the alignment between the corresponding feature vectors in $\mathbf{P}$ and $\mathbf{Z}$, as follows:
@@
4 Experiments
Datasets and baselines: We extensively evaluate our approach on 13 diverse datasets: ImageNet [42], Caltech101 [43], DTD [44], EuroSAT [18], FGVCAircraft [45], Food101 [46], Flowers102 [47], OxfordPets [48], SUN397 [49], StandfordCars [50], CIFAR10\100 [51], and UCF101 [52]. We perform comparative analysis with four state-of-the-art (SOTA) unsupervised adaptation methods: CLIP [1], UPL [14], POUF [13], and LaFTer [12].
Implementation Details: For all experiments, unless otherwise specified, we utilize a ViT/B-32 CLIP pre-trained by OpenAI [1]. In the context of unsupervised fine-tuning, we specifically target the layer-normalization weights of the image encoder and the textual prototypes. This strategy is effective and stable for adapting models with noisy supervision [53, 54]. The text encoder of CLIP will be excluded after constructing the textual prototypes. Input images are standardized to a size of 224 × 224. During training, we apply a series of data augmentation techniques, including RandomResizedCrop, Flip, RandAugment [55] as a strong augmentation and Resize+RandomCrop for generating pseudo-labels. For testing, we extract a Center crop after resizing the image to $224$. We utilize AdamW optimizer [56] with a cosine learning rate schedule. For fair comparisons, we reproduce the results of UPL, POUF, and LaFTer using their publicly available code. For more implementation details, please refer to the supplementary materials.
Results: Table 1 reports the top-1 accuracy of our proposed method DPA and four SOTA baselines: CLIP, UPL, POUF, and LaFTer on 13 datasets. Our approach consistently outperforms zero-shot CLIP, improving top-1 accuracy by $+7.83\%$ on average. Compared to UPL, which optimizes prompts, DPA still achieves consistent improvements with an average gain of $+8.24\%$ with minimal additional computational overhead despite slightly larger parameters. DPA also surpasses POUF and LaFTer, achieving average increases of $+7.47\%$ and $+3.93\%$, respectively, across the 13 datasets. Notably, DPA outperforms LaFTer without requiring additional text corpora or further pre-training.
Table 1: Comparison with SOTA unsupervised adaption methods.

### v2

@@ Abstract
Vision-language models (VLMs), e.g., CLIP, have shown remarkable potential in zero-shot image classification. However, adapting these models to new domains remains challenging, especially in unsupervised settings where labeled data is unavailable. Recent research has proposed pseudo-labeling approaches to adapt CLIP in an unsupervised manner using unlabeled target data. Nonetheless, these methods struggle due to noisy pseudo-labels resulting from the misalignment between CLIP’s visual and textual representations. This study introduces DPA, an unsupervised domain adaptation method for VLMs. DPA introduces the concept of dual prototypes, acting as distinct classifiers, along with the convex combination of their outputs, thereby leading to accurate pseudo-label construction. Next, it ranks pseudo-labels to facilitate robust self-training, particularly during early training. Finally, it addresses visual-textual misalignment by aligning textual prototypes with image prototypes to further improve the adaptation performance. Experiments on 13 downstream vision tasks demonstrate that DPA significantly outperforms zero-shot CLIP and the state-of-the-art unsupervised adaptation baselines.
@@ 1 Introduction
Context and background: Vision-language models (VLMs) [38, 20, 56, 27] have shown promising potential in zero-shot image classification. One notable example is CLIP [38]. Despite the impressive zero-shot capabilities of CLIP, its performance can be impacted by the discrepancy between the pretraining image-text pairs and the downstream task images [38, 1]. To address this limitation, several studies have attempted to enhance CLIP transfer performance on downstream tasks by leveraging limited labeled samples from the target domains [59, 61, 60, 21, 22]. However, in many practical applications, such as security and medical diagnostics, collecting labeled samples can be particularly challenging due to high labeling costs and data privacy concerns. In such contexts, unsupervised learning presents a promising alternative.
Several contributions for adapting CLIP to a target domain using an unlabeled dataset have recently been introduced [25, 33, 45, 18, 37, 17]. A common approach involves leveraging CLIP to generate pseudo-labels, which are then used to fine-tune CLIP in an unsupervised manner. However, this method encounters significant challenges due to the noisy pseudo-labels generated by CLIP. Despite efforts by recent methods for the unsupervised adaptation of CLIP, a significant modality gap persists between the text and vision representations [17, 29], leading to inaccurate pseudo-labels causing confirmation bias during adaptation.
Motivation: Recent studies highlight a significant factor behind the performance issues of CLIP in unsupervised adaptation scenarios: the visual domain gap between the source images used to train CLIP and the target images typically occurs when the target samples originate from an uncommon domain [17]. As shown in Fig. 1-(a), the t-SNE projection of visual and textual embeddings on the EuroSAT dataset [12] reveals a significant misalignment, resulting in misclassifications. Contemporary methods (e.g., [17]) attempt to address this misalignment by learning a projection space to mitigate the issue and using label propagation [19] to improve pseudo-labels. However, they could be limited as they are either computationally expensive and/or show suboptimal performance in an inductive setting. Further, on datasets with a large number of classes, they struggle to perform well, necessitating turning off label propagation and using model predictions. Effectively addressing the challenges of unsupervised adaptation of CLIP in an inductive setting is crucial for improving CLIP’s performance in unsupervised adaptation scenarios.
@@
Our proposal: To address these challenges, we propose DPA, a novel unsupervised adaptation method for VLMs (Figure 2). DPA aims at addressing the domain gap between visual and textual representations in downstream tasks. DPA introduces the idea of dual prototypes, namely image and textual prototypes, which act as distinct classifiers, and their outputs are fused via a convex combination towards generating accurate pseudo-labels. There are two main reasons for introducing dual prototypes. Firstly, the misalignment between the image representation and its textual representation in zero-shot CLIP, due to domain shift, often leads to inaccurate pseudo-labels (Figure 1-(a)). Secondly, image prototypes, which tend to be less affected by noise, as shown in Fig. 1, are generally closer to the true image representation than textual prototypes. Additionally, we tackle the challenge of misalignment between visual and textual embeddings by aligning textual prototypes with image prototypes. This alignment process further enhances the performance of unsupervised CLIP adaptation.
Contributions: 1) We propose a novel dual-modality prototypic alignment framework, namely DPA for adapting VLMs in an unsupervised manner. Our approach introduces a dual-prototype pseudo-labeling strategy, where the visual prototypes serve as non-parametric classifiers to bring target visual embeddings closer to their respective class prototypes. The learnable textual prototypes, initialized by ensembling different prompt templates using zero-shot CLIP, provide rich task-specific semantics that CLIP visual embeddings typically lack. 2) We propose ranking pseudo-labels in the classification loss to mitigate noise, particularly during early training. Furthermore, we address visual-textual misalignment by aligning textual prototypes with image prototypes, ensuring the separation of different semantics while aligning with relevant ones. 3) Extensive experiments on 13 downstream vision tasks demonstrate consistent and significant performance enhancements over zero-shot CLIP and the state-of-the-art baselines.
@@ 2 Related Work
Large-scale Vision-Language (VL) Models: Among other vision-language models [8, 41, 20, 38, 6, 56, 16], CLIP [38] stands out as a pioneering example, aligning visual and textual features via a contrastive objective on a vast web-crawled collection of image-text pairs. This alignment empowers CLIP to generalize to diverse downstream classification tasks effectively. Building on CLIP’s success, subsequent research has explored methods for efficiently transferring the pre-trained model to handle diverse downstream tasks with limited labeled target data. These approaches often leverage vision-specific adapters [10, 59, 58, 63], language-specific adapters [61, 60, 22], or both [21]. However, these techniques require a minimum number of labeled samples, which poses challenges due to data scarcity, annotation expenses, privacy issues, and practical limitations. Recent advancements have seen the emergence of unsupervised adaptation techniques, focusing on tailoring CLIP to target tasks using unlabeled datasets [18, 45, 17, 33, 25, 37]. However, these methods still depend on additional supervision signals, such as fine-tuning classifiers with a large language model such as GPT-3, as seen in LaFTer [33]. They also often require significant computational resources, as exemplified by MUST [25]. ReCLIP [17] tackles misaligned embeddings through source-free domain learning, utilizing a projection space to align these embeddings and employing pseudo-labels for self-training. Despite progress, there remains significant potential to improve the performance of unsupervised adaptation techniques across various image classification benchmarks. In contrast, our work introduces a fully label-free approach for adapting VLMs to a target task by generating more accurate pseudo-labels grounded on image and text prototypes.
Prototype-based Learning: Prototype-based learning uses a fixed set of distinctive prototypes representing the data, and then the learning is achieved by comparing the test samples directly with these prototypes. Prototype-based learning has been extensively studied in various contexts, including few-shot classification [42, 28], unsupervised learning [50, 54] and supervised classification [55, 32]. Our method innovatively incorporates two types of prototypes, image prototypes, and textual prototypes, to harness both prototypes to ultimately generate accurate pseudo-labels and effectively adapt CLIP to the target dataset.
Pseudo-labeling: Pseudo-labeling is popular in various domains, including vision [57, 40] and vision-language tasks [25, 33, 18, 17]. The pseudo-labeling pipeline of DPA leverages consistency regularization [43] to produce consistent predictions across different data augmentations, thereby enhancing pseudo-labeling accuracy.
@@ 3 Methodology
This study addresses the task of unsupervised adaptation of vision-language models (i.e., CLIP) for image classification. Pre-trained CLIP consists of visual encoder $E_{v}$ and textual encoder $E_{t}$. The target dataset $\mathcal{D}_{t}=\left\{\mathcal{X}_{t}\right\}$ consists of unlabeled images $\left\{x_{i}\right\}_{i=1}^{N}$, where $x_{i}\in\mathcal{X}_{t}$, and we have unique class names $\mathcal{C}$ =$=\{c_{j}\}_{j=1}^{C}$ for the unlabeled target dataset. We use a pre-trained CLIP to process the unlabeled target data $\mathcal{D}_{t}$. Initially, the source model assigns a pseudo-label to each unlabeled target image $x_{i}$. During the adaptation phase, we employ two prototypes to generate pseudo-labels for adapting CLIP to the target domain.
@@ 3.1 Zero-shot Visual Classification in CLIP
where $f_{i}^{s}=E_{v}(x_{i}^{s})$ and $z_{i}=E_{t}(\mathbf{t}^{s}_{i})$, with both $f_{i}^{s}\in\mathbb{R}^{1\times d}$ and $z_{i}\in\mathbb{R}^{1\times d}$. The parameter $\tau$ represents a learned temperature, while $m$ indicates the mini-batch size. For inference, given an unlabeled target image $x$ and a list of class names $\mathcal{C}$, we leverage the pre-trained textual encoder $E_{t}$ to generate text embeddings for each class. Specifically, a series of well-crafted prompts are fed into $E_{t}$, producing a set of text embeddings for each class, denoted as $\mathbf{Z}$ where $\mathbf{Z}\in\mathbb{R}^{C\times d}$. The zero-shot prediction $\hat{y}$ is then obtained as follows:
$\hat{y}=\arg\max_{c}(f\cdot\mathbf{Z}^{T})$ (2)
@@ 3.2 Proposed Framework (DPA)
Motivation: While zero-shot CLIP shows impressive performance across diverse domains, it still exhibits notable limitations compared to models adapted under supervised conditions due to the misalignment between image and corresponding textual representations [38]. As depicted in Fig. 1-(a), this discrepancy is evident in CLIP when applied to uncommon domains. Such disparities can lead to inaccurate pseudo-labels, especially when utilizing zero-shot CLIP to generate pseudo-labels for the unsupervised adaptation [17]. Our objective is to adapt CLIP to the target domain using its unlabeled samples by bridging the gap between image representations and their textual representations, all without dependence on ground-truth labels. Simultaneously, we seek to achieve efficient parameter-based unsupervised adaptation for CLIP. Our objective relies on two key components: (1) ensuring the accuracy of the pseudo-labels generated for the unlabeled samples and (2) aligning textual representations with their corresponding visual representations. To generate the pseudo-labels for adapting CLIP to the target domain, we introduce dual prototypes that act as two different classifiers and fuse their complementary outputs via a convex combination. Furthermore, we propose visual-textual prototype alignment towards further improving the adaptation performance.
@@
Image Prototypes Construction: To construct the image prototypes, CLIP’s image encoder processes a weakly-augmented unlabeled image $\alpha(x)$ and generates a feature representation $f=E_{v}(\alpha(x))$, where $f\in\mathbb{R}^{1\times d}$. Without class labels, we utilize Eq. 1 and Eq. 2 to generate pseudo-labels $\hat{y}$ from zero-shot CLIP. To obtain the image prototype $\mathbf{P}_{j}$ for a class $c_{j}$, we average the feature representation $f$ of the unlabeled images assigned to class $c_{j}$ as follows:
$\mathbf{P}_{j}=\frac{1}{N_{j}}\sum_{i=1}^{N}\!\mathbb{I}\{\hat{y_{i}}=j\!\}\!\;f_{i}$ (4)
where $N_{j}=\sum_{i=1}^{N}\!\mathbb{I}\{\hat{y_{i}}=j\!\}\!$. For the set of $C$ classes, we define the image prototypes as $\mathbf{P}\in\mathbb{R}^{C\times d}$—the image prototypes function as a non-parametric classifier. Given that the pseudo-labels generated by zero-shot CLIP are initially noisy, our approach, inspired by previous works [62, 49, 53], employs a memory bank $\mathbf{MB}$. This memory bank facilitates the incremental update of image prototypes in a non-parametric way throughout the training process. Specifically, $\mathbf{MB}$ stores the feature representations of all unlabeled target samples along with their corresponding pseudo-labels: $\mathbf{MB}=\{(f_{i},\hat{y}_{i})\}_{i=1}^{N}$. As the image prototypes are non-parametric, we iteratively refine them using the memory bank as follows: (1) We initially calculate the image prototypes using Eq. 4, utilizing the image features and the pseudo-labels generated from zero-shot CLIP across the entire target dataset. (2) During training, we continuously update the memory bank by storing the image representations alongside their corresponding pseudo-labels. (3) At the end of each epoch, we update the prototypes $\mathbf{P}_{j}$ using Eq. 4, incorporating the image representations and their corresponding pseudo-labels from $\mathbf{MB}$. (4) We repeat steps (2) and (3) until the completion of the training process. DPA maximizes the utilization of all available training samples to enhance the image prototypes. This iterative process ensures that each prototype continuously evolves, progressively assimilating relevant knowledge from individual training samples throughout the model training phase while mitigating the error propagation commonly associated with pseudo-labels.
Pseudo-labels Generation: Both image and textual prototypes serve as distinct classifiers for the unlabeled target dataset. We generate a pseudo-label (PL) for the weakly-augmented unlabeled target image $\alpha(x)$ from these two prototypes. Initially, we determine the similarity between the image features and the textual prototypes as follows:
@@
Following [26], we perform distribution alignment (DA) to prevent the model’s prediction from collapsing onto specific classes. Here, $DA(p_{t})=p_{t}/\bar{p}_{t}$, where $\bar{p}_{t}$ represents a running average of $p_{t}$ during training, and $\beta$ is a hyperparameter. Finally, the PL is obtained as $\hat{y}=\arg\max_{c}\hat{p}$. This process of generating pseudo-labels improves the accuracy of the PL during training, indicating that these PLs may initially be noisy (see Fig. 4). Therefore, assigning uniform weights to all pseudo-labels could hinder the adaptation process. To address this challenge, we propose adjusting the classification loss weight for the pseudo-label $\hat{y}$ based on its similarity to the textual and image prototypes, as follows:
@@
$\mathcal{L}_{st}=-\;\mathbb{E}_{x\in\mathcal{X}_{t}}\;\left[w_{x}\cdot\sum_{j=1}^{C}\!\mathbb{I}\{\hat{y}=j\!\}\!\;log\left(p_{\mathcal{A}(x)}\right)\right],$ (9)
where $w_{x}$ represents the weight calculated from Eq. 8, and $p_{\mathcal{A}(x)}=E_{v}(\mathcal{A}(x))\cdot\mathbf{Z}^{T}$ represents the probabilistic output for the strongly-augmented image $\mathcal{A}(x)$. To alleviate the confirmation bias induced by CLIP [48], we adopt “fairness” regularization as suggested in [25] as follows:
$\mathcal{L}_{reg}=-\frac{1}{C}\sum_{j=1}^{C}\log\left(\bar{p}_{\mathcal{A}(x_{j})}\right),$ (10)
where $\bar{p}_{\mathcal{A}(x)}$ represents the model’s average prediction from the strongly augmented images across the batch. By incorporating this fairness regularization, we aim to encourage the model to make more uniform predictions across classes, reducing the tendency to overfit the PLs and promoting a more balanced adaptation to the target domain. Eq. 9 and Eq. 10 are used to update both the image encoder and the textual prototypes.
Prototypes Alignment: During training, the textual prototypes undergo continual updates to align with the associated image representation, leveraging pseudo-labels using Eq. 9 and Eq. 10. Conversely, the refinement of the image prototypes occurs through averaging image features, as detailed in Section 3.2. We exploit the image prototypes to provide additional supervision for updating textual prototypes beyond pseudo-label alone. These image prototypes, less influenced by erroneous pseudo-labels and closer to image features (as illustrated in Fig. 1), offer extra guidance for refining textual prototypes. Our primary objective is to optimize the relationship between image prototypes $\mathbf{P}$ and textual prototypes $\mathbf{Z}$ by maximizing the (cosine) similarity between these feature sets. To accomplish this optimization, we employ the InfoNCE loss [35] to directly assess the alignment between the corresponding feature vectors in $\mathbf{P}$ and $\mathbf{Z}$, as follows:
@@
Table 1: Comparison of state-of-the-Art unsupervised adaptation methods using the ViT-B/32 backbone.
4 Experiments
Datasets and Baselines: We extensively evaluate on 13 diverse datasets: ImageNet [7], Caltech101 [9], DTD [4], EuroSAT [12], FGVCAircraft [31], Food101 [3], Flowers102 [34], OxfordPets [36], SUN397 [52], StandfordCars [23], CIFAR10\100 [24], and UCF101 [44]. We perform comparative analysis with four state-of-the-art (SOTA) unsupervised adaptation methods: CLIP [38], UPL [18], POUF [45], and LaFTer [33].
Implementation Details: For all experiments, unless otherwise specified, we utilize a ViT/B-32 CLIP pre-trained by OpenAI [38]. In the context of unsupervised fine-tuning, we specifically target the layer-normalization weights of the image encoder and the textual prototypes. This strategy is effective and stable for adapting models with noisy supervision [2, 46]. The text encoder of CLIP will be excluded after constructing the textual prototypes. Input images are standardized to a size of 224 × 224. During training, we apply a series of data augmentation techniques, including RandomResizedCrop, Flip, RandAugment [5] as a strong augmentation and Resize+RandomCrop for generating pseudo-labels. For testing, we extract a Center crop after resizing the image to $224$. We utilize AdamW optimizer [30] with a cosine learning rate schedule. To ensure fair comparisons, we reproduced the results of UPL, POUF, and LaFTer using their publicly available code. Most state-of-the-art (SOTA) methods use different backbones, and UPL’s reported results were only for 16 shots, unlike our approach, which provides results for the entire training dataset. For additional details on the implementation, see the supplementary materials.
Results: Table 1 shows that our approach consistently outperforms zero-shot CLIP, improving top-1 accuracy by $+7.83\%$ on average. Compared to UPL, which optimizes prompts, DPA still achieves consistent improvements with an average gain of $+8.24\%$ with minimal additional computational overhead despite slightly larger parameters. We observe that UPL’s performance on the entire unlabeled training dataset is inferior to that of zero-shot CLIP. This discrepancy may stem from the fact that UPL generates offline pseudo-labels using zero-shot CLIP, which can lack confidence, particularly when dealing with fine-grained classes in certain datasets [33]. DPA also surpasses POUF and LaFTer, achieving average increases of $+7.47\%$ and $+3.93\%$, respectively, across the 13 datasets without requiring additional text corpora or further pre-training. We note significantly larger performance improvement on the ESAT dataset compared to other datasets. We believe this is because zero-shot CLIP exhibits limitations when applied to specialized, complex, or abstract tasks, such as satellite image classification [38]. This suggests there is potential for improvement through additional supervision. This behavior aligns with findings from previous studies. LaFTer achieved a remarkable +26.12% improvement over zero-shot CLIP. ReCLIP reported gains exceeding +26.96% (Table 8). The 10% improvement by our method (DPA) over LaFTer highlights the effectiveness of our pseudo-labeling approach.